# MAYA: Addressing Inconsistencies in Generative Password Guessing through a Unified Benchmark

### v2

@@ 5.3 RQ3 - How Sensitive are Models to Training Dataset Size?
We assess the ability of generative models to capture the full password distribution when trained on different data subset sizes. Models were trained on up to seven different subset sizes across four datasets, with a minimum initial size of 1M passwords. Results in Table IV illustrate the models’ varying performance. Transformer-based models (PassGPT and VGPT2) consistently improve as the training dataset size increases, with VGPT2, in particular, struggling on smaller subsets. FLA also generally benefits from larger training subsets, although the rate of improvement tapers off between $63e$ 6 and $71e$ 7 passwords. We also observe an anomalous behavior on LinkedIn, where performance drops significantly after $71e$ 7—or $25\%$ of the dataset size. PLR-GAN excels on small training subsets, but its performance declines around $30\%$ of the dataset size before improving again, ultimately achieving its best performance on the full dataset. PassFlow remains the most consistent model across subset sizes, with minimal performance variation. In contrast, PassGAN exhibits highly variable results, with performance peaks at different dataset sizes, making its behavior less predictable. Overall, transformer models show the most significant improvement with increasing training data, whereas other architectures tend to exhibit flat or minimal gains, challenging the conventional assumption that more training data universally leads to better performance.
@@ 5.4.1 Cross-community
As shown in Table V, Cross-community generalization is strongly model- and dataset-dependent. PassGAN, PLR-GAN, PassFlow, and VGPT2 achieve their highest performance on LinkedIn when trained on RockYou, suggesting that the distribution learned from RockYou generalizes well despite community differences. However, stronger models such as PassGPT and FLA experience notable performance drops in the same scenario. This contrast leads us to hypothesize that the apparent generalization success of weaker models stems from their limited capacity to capture fine-grained details of the training distribution—resulting in broader, but less accurate, generalization. We also note how all approaches struggle to generalize when training on 000webhost, highlighting that performance significantly degrades when the source and target distributions diverge too strongly (see Table X).
@@ 5.4.2 Cross-culture
In the cross-culture scenario, model performance follows a more expected pattern, as illustrated in Table VI: models trained on one cultural or linguistic context tend to perform significantly worse when evaluated on datasets from a different one. This highlights the challenges of generalizing across culturally distinct password distributions and underscores the importance of training data alignment with the target population. However, some interesting behaviors still emerge. Despite the challenging setting, generative models still manage to guess a non-negligible portion of the target passwords. Notably, models that typically underperform in other settings—such as PassFlow—can surpass stronger models like PassGPT when there is a significant mismatch between source and target distributions. This suggests that certain models may possess greater flexibility or robustness in the face of distributional shifts, even if they are less effective under ideal, in-distribution conditions.
@@ 5.5.1 Analysis by Password Frequency
From each test dataset, we created three subsets based on password frequency: top $5\%$, top $10\%$, and bottom $90\%$. Since nearly $90\%$ of test passwords across all datasets are unique (see Figure 1), these subsets allow us to evaluate models under three scenarios: (1) very common, (2) common, and (3) rare passwords. We excluded LinkedIn and Ashley Madison, as they consist entirely of unique passwords. Results, shown in Figure 4, are reported as weighted averages. As expected, all models perform better on common passwords. While the drop from the top $5\%$ to the top $10\%$ is small, performance declines sharply when focusing on the bottom $90\%$. Still, models are surprisingly capable of guessing rare passwords, with PassGPT and FLA guessing close to $40\%$ and $50\%$, respectively. In Figure 5, we leverage PassFlow’s encoder to visualize password embeddings and observe that the bottom $90\%$ of passwords tend to form clusters around the top $10\%$. Since PassFlow’s latent space is smooth [42], this behavior suggests that many rare passwords are slight variations of frequently used ones, which may explain why models are still able to guess a non-negligible portion of them despite their low frequency.
Figure 4: Guessing performance by password frequency group.
Figure 5: t-SNE plot showing the projection of Top 10% and Bottom 90% passwords in PassFlow’s latent space. Embedding computed using PassFlow’s encoder.
TABLE VII: Guessing performance by password pattern. Values expressed as percentage of guessed passwords for each pattern.
Model r1 r2 r3 r4 r5 r6 r7 r8 r9 r10 r11 r12 r13 r14 r15 r16 r17 r18 r19 PassGAN 6.94 7.65 0.16 28.31 0.30 3.55 0.10 0.23 0.08 4.12 0.10 0.53 0.12 24.17 0.10 0.01 0.07 0.06 7.53 PLR-GAN 12.60 13.74 1.35 39.95 0.67 7.31 0.77 0.74 0.27 8.39 0.64 2.14 0.47 34.17 0.35 0.03 0.15 1.02 12.49 PassFlow 10.38 11.35 1.41 27.92 5.34 4.82 0.70 0.30 0.05 5.01 0.41 2.29 0.21 24.02 1.79 0.45 0.16 0.39 9.54 PassGPT 38.47 39.66 33.74 46.55 25.82 29.81 20.04 19.90 11.19 33.43 17.88 29.62 10.39 41.68 18.19 12.72 5.04 21.63 36.16 VGPT2 12.09 13.11 2.38 29.39 2.59 5.37 1.22 0.90 0.33 6.10 0.95 2.15 0.53 25.19 0.61 0.04 0.09 1.07 9.74 FLA 46.51 48.14 38.49 67.56 17.51 39.38 21.02 26.28 14.20 44.63 21.32 26.39 11.88 59.87 13.92 8.50 4.58 24.57 48.09
@@ 5.5.2 Analysis by Password Length
We analyzed matches by password length, ranging from 4 to 12 characters. Results, computed using a weighted average across all datasets, are shown in Figure 6. Notably, all models perform well on short passwords, with FLA nearly achieving a $100\%$ match rate. However, performance declines as length increases. Only FLA and PassGPT maintain a substantial guessing rate beyond 8 characters. Notably, consistent with findings from RQ1, PassGPT surpasses FLA from length 11 onward. PassFlow initially performs on par with PassGPT for shorter lengths but undergoes a sharp decline, becoming the weakest performer from length 8 onward.
Figure 6: Guessing performance by password length. Dots represent the percentage of generated passwords for each length.
@@ 5.5.3 Analysis by Password Patterns
We evaluated the models’ ability to guess passwords based on the structural patterns defined in Table III. The results, presented in Table VII as a weighted average across all datasets, show that PassGPT and FLA consistently achieve the strongest performance, standing out as the only models capable of successfully guessing passwords across all defined patterns. In contrast, the remaining models show limited pattern coverage, particularly struggling with complex password structures that include special characters (r5, r7, r8) or uncommon character combinations (r11, r12, r13). GAN-based models fare especially poorly on passwords composed entirely of special characters (r5), significantly underperforming PassFlow and VGPT2. This suggests a clear limitation in capturing rare or unconventional character sets. In contrast, all models exhibit strong performance on digit-only passwords (r4), indicating that simple patterns are consistently learned regardless of the architecture.
@@
We assess the extent to which generative models’ learned distributions overlap and investigate the effectiveness of combining multiple models to enhance guessing performance.
@@
We investigate the first aspect of this RQ utilizing two primary metrics: the Jaccard Index and the Mergeability Index. Their mathematical definitions are provided in Appendix B, in Equations 1 and 2, respectively. The Jaccard Index provides a quantitative measure of the overlap between two sets of generated passwords, with a value closer to 1 indicating that the two models produce similar distributions, and a value closer to 0 indicating that they generate distinct sets of passwords. As illustrated in Figure 7(a), the FLA-PassGPT pair is the only combination with a Jaccard Index greater than $0.1$, suggesting that these two models share some similarities in their password generation patterns. All other pairs exhibit much lower Jaccard values. Interestingly, despite PassGAN and PLR-GAN both being based on GANs, their Jaccard Index is quite low. Similarly, VGPT2-PassGPT also shows a low Jaccard Index. These results highlight that, even within the same architectural family, models may produce highly distinct password distributions, suggesting that combining multiple models could be beneficial for maximizing coverage of the password space.
@@
While the Jaccard Index provides insights into the overlap of generated passwords, it does not account for their effectiveness in terms of actual matches with real-world data. To address this, we introduce the Mergeability Index, a complementary metric that measures the benefits of combining the output of two models based on their successful guesses. It captures the performance improvement achieved by merging the guesses relative to the best-performing model. A Mergeability Index close to 0 indicates that the models guess mostly the same set of passwords, while a value close to 1 indicates largely distinct matching password sets. The results are shown in Figure 7(b). Interestingly, FLA exhibits the lowest Mergeability across all models, indicating that its guessed passwords are generally a superset of those generated by the other models. This suggests that FLA has already captured a significant portion of the password space, and merging it with other models provides only limited additional benefit. In contrast, PassFlow, PassGAN, PLR-GAN, and VGPT2 show high Mergeability, with their varying combinations improving successful guesses by 18% to 30%. This suggests that these models guess highly distinct sets of passwords, meaning they are each capturing different aspects of the password space.
@@
We assess the potential of a multi-model attack by evaluating the performance gain achieved through the combination of the generated password sets from $n$ models. Our results, presented in Figure 8, show the relative gain with respect to the best-performing single model, computed as percentage points on the test set. We identify the following optimal model combinations for different values of $n$: $n_{1}=\{FLA\}$, $n_{2}=n_{1}+PassGPT$, $n_{3}=n_{2}+PLR$-$GAN$, $n_{4}=n_{3}+PassFlow$, $n_{5}=n_{4}+PassGAN$, and $n_{6}=n_{5}+VGPT2$. In datasets like LinkedIn and RockYou, combining models yields significant gains over FLA alone—especially PassGPT. Remarkably, combining PassGPT and FLA provides a larger average improvement than adding the remaining four models combined. These results mirror those presented in Figure 7(b), indicating that FLA and PassGPT capture somewhat complementary parts of the password space. However, in datasets such as 000webhost and Mailru, FLA alone performs nearly as well as the combination of all models, with only a $1$ percentage point improvement in the multi-model attack.
@@ 5.7.1 Metrics-Based Evaluation
IMD is an intrinsic, multi-scale metric that compares the data manifolds of real and generated samples and is applicable across diverse domains. However, its effectiveness is highly sensitive to the choice of feature representation, which can introduce bias into the distance estimation.
@@
TABLE VIII: Distance between human- and generative model-created passwords.
@@
FLA 12% -15% -1% 31% 172% 0%
To contextualize the metric values, we define two baselines: (1) a soft lower bound (optimal case), defined as the metric value between test and training passwords; (2) a soft upper bound (worst case), defined as the metric value between randomly generated and test passwords. For each model, we normalize the metric value between lower ($0\%$) and upper ($100\%$) bounds, as displayed in Table VIII. Values close to zero indicate closer alignment between the generated and human-created password distributions. GAN-based models demonstrate generally good performance across all metrics, with PLR consistently achieving lower values than PassGAN. PassGPT generates passwords that closely resemble human-created ones, likely due to its underlying GPT2 architecture, which is well-suited for modeling textual distributions. In contrast, PassFlow demonstrates the weakest performance, particularly in the IMD metric, where its score exceeds the soft upper bound by a factor of two. We attribute this behavior to PassFlow’s Gaussian Smoothing (GS) strategy, which introduces random perturbations in the generation process. While GS enhances uniqueness [42], it artificially distorts the distribution of generated passwords, ultimately reducing their quality. VGPT2 and FLA exhibit similar trends, performing strongly across most metrics except for IMD, where elevated values across most models suggest that they all struggle to fully capture certain characteristics of human passwords. Conversely, MTopDiv consistently yields low values, indicating that the models effectively avoid common mode failures, such as dropping, collapse, and invention.
@@ 5.7.2 Length Distribution in Generated Passwords
Figure 9 compares the length distribution of generated passwords with that of real passwords. Among the models, PassFlow, FLA, and VGPT2 show the largest divergence from the real distribution: PassFlow tends to generate shorter passwords, FLA overestimates the probability of 8-character passwords, and VGPT2 of longer ones. In contrast, PassGAN, PLR-GAN, and PassGPT produce distributions that more closely align with real data, reinforcing previous observations.
@@ 5.7.3 Why Does IMD Yield High Values?
Figure 9 reveals a clear correlation between length distribution and IMD values. PassFlow, which exhibits the largest deviation from the real distribution, also records the highest IMD score. A similar trend is observed for FLA and VGPT2. Conversely, models such as PassGAN, PLR-GAN, and PassGPT, which more closely replicate the real-world length distribution, achieve lower IMD scores. Interestingly, Table VIII shows that models such as PassFlow and FLA yield higher IMD scores than those associated with the random-password soft upper bound. We posit that this counterintuitive result stems from the uniform length distribution of random passwords. Although this distribution diverges significantly from that of real passwords, it nonetheless includes longer passwords that PassFlow and FLA struggle to generate. These observations suggest that IMD is highly sensitive to length distribution, favoring models that replicate it more accurately. For an extended analysis of IMD, we refer to Appendix B.
TABLE IX: Comparison summary of the selected approaches.
Metric PassGAN PLR-GAN PassFlow PassGPT VGPT2 FLA
@@
   Scenario: In-Distribution $\uparrow$ 8.85 14.88 12.76 37.21 11.79 49.21
   Scenario: Cross-Community $\uparrow$ 3.62 6.16 7.13 17.72 4.76 27.72
   Scenario: Cross-Culture $\uparrow$ 6.52 9.70 16.17 14.24 7.01 23.27
@@
   Frequency: Common (Top 5%) $\uparrow$ 27.87 36.86 31.20 68.20 32.83 83.37
   Frequency: Common (Top 10%) $\uparrow$ 22.73 30.63 25.73 63.45 27.23 76.60
   Frequency: Rare (Bottom 90%) $\uparrow$ 8.75 13.61 12.28 34.51 10.83 47.03
@@
   Length: Short (4-7 Chars) $\uparrow$ 19.91 32.96 28.16 55.62 29.65 72.97
   Length: Medium (8-10 Chars) $\uparrow$ 2.48 4.47 1.33 26.93 2.16 36.38
   Length: Long (11-12 Chars) $\uparrow$ 0.07 0.24 0.00 12.58 0.05 11.6
@@
   Pattern: Simple (1 Char Class) $\uparrow$ 14.06 21.71 16.22 41.16 17.85 53.52
   Pattern: Moderate (2 Char Classes) $\uparrow$ 3.46 7.15 4.71 29.56 5.26 38.94
   Pattern: Complex (3 Char Classes) $\uparrow$ 0.08 0.27 0.05 11.19 0.33 14.20
@@
   Train Set Size Sensitivity (%) $\downarrow$ 12.56 3.54 1.90 13.16 35.49 7.58
   Cross-Community Loss (%) $\downarrow$ 24.16 32.42 6.84 42.15 32.49 30.13
   Cross-Culture Loss (%) $\downarrow$ 48.67 49.27 14.58 66.15 57.11 58.53
@@
   Uniqueness (%) $\uparrow$ 54.59 71.13 90.10 73.22 91.41 100.00
   Humanness Distance (%) $\downarrow$ 19.83 4.80 70.16 2.40 44.00 38.50

### v4

@@ 5.3 RQ3 - How Sensitive are Models to Training Dataset Size?
We assess the ability of the models to capture the full password distribution when trained on different data subset sizes. Models were trained on up to seven different subset sizes across four datasets, with a minimum initial size of 1M passwords. Results in Table VI illustrate the models’ varying performance. PCFG and both transformer-based models (PassGPT and VGPT2) consistently improve as the training dataset size increases, with VGPT2, in particular, struggling on smaller subsets. FLA also generally benefits from larger training subsets, although the rate of improvement tapers off between $3e6$ and $1e7$ passwords. We also observe an anomalous behavior on LinkedIn, where performance drops significantly after 1e71e7—or $25\%$ of the dataset size. PLR-GAN excels on small training subsets, but its performance declines around $30\%$ of the dataset size before improving again, ultimately achieving its best performance on the full dataset. OMEN and PassFlow remain the most consistent models across subset sizes, with minimal performance variation. In contrast, PassGAN exhibits highly variable results, with performance peaks at different dataset sizes, making its behavior less predictable.
Overall, transformer models show the most significant improvement with increasing training data, whereas other architectures tend to exhibit flat or minimal gains, challenging the conventional assumption that more training data universally leads to better performance.
@@ 5.4.1 Cross-community
As shown in Table VII, Cross-community generalization is strongly model- and dataset-dependent. OMEN, PassGAN, PLR-GAN, PassFlow, and VGPT2 achieve their highest performance on LinkedIn when trained on RockYou, suggesting that the distribution learned from RockYou generalizes well despite community differences. However, stronger models such as PassGPT, PCFG, and FLA experience notable performance drops in the same scenario. This contrast leads us to hypothesize that the apparent generalization success of weaker models stems from their limited capacity to capture fine-grained details of the training distribution—resulting in broader, but less accurate, generalization. We also note how all approaches struggle to generalize when training on 000webhost, highlighting that performance significantly degrades when the source and target distributions diverge too strongly (see Table XIII).
@@ 5.4.2 Cross-culture
In the cross-culture scenario, model performance follows a more expected pattern, as illustrated in Table VIII: models trained on one cultural or linguistic context tend to perform significantly worse when evaluated on datasets from a different one. This highlights the challenges of generalizing across culturally distinct password distributions and underscores the importance of training data alignment with the target population. However, some interesting behaviors still emerge. Despite the challenging setting, models still manage to guess a non-negligible portion of the target passwords. Notably, generative models that typically underperform in other settings—such as PassFlow—can surpass stronger models like PassGPT, OMEN, and PCFG, when there is a significant mismatch between source and target distributions. This suggests that certain models may possess greater flexibility or robustness in the face of distributional shifts, even if they are less effective under ideal, in-distribution conditions.
Figure 4: Guessing performance by password frequency.
@@ 5.5.1 Analysis by Password Frequency
From each test dataset, we created three subsets based on password frequency: top $5\%$, top $10\%$, and bottom $90\%$. Since nearly $90\%$ of test passwords across all datasets are unique (see Figure 1), these subsets allow us to evaluate models under three scenarios: (1) very common, (2) common, and (3) rare passwords. We excluded LinkedIn and Ashley Madison, as they consist nearly entirely of unique passwords to avoid biased findings. Results, shown in Figure 4, are reported as weighted averages. As expected, all models perform better on common passwords. While the drop from the top $5\%$ to the top $10\%$ is small, performance declines sharply when focusing on the bottom $90\%$. Still, models are surprisingly capable of guessing rare passwords, with PassGPT and PCFG guessing close to $40\%$, and FLA close to $50\%$. In Figure 5, we leverage PassFlow’s encoder to visualize password embeddings and observe that the bottom $90\%$ of passwords tend to form clusters around the top $10\%$. Since PassFlow’s latent space is smooth [50], this behavior suggests that many rare passwords are slight variations of frequently used ones, which may explain why models are still able to guess a non-negligible portion of them despite their low frequency.
Figure 5: t-SNE plot showing the projection of Top 10% and Bottom 90% passwords in PassFlow’s latent space.
TABLE IX: Guessing performance by password pattern. Values expressed as percentage of guessed passwords for each pattern.
Model r1 r2 r3 r4 r5 r6 r7 r8 r9 r10 r11 r12 r13 r14 r15 r16 r17 r18 r19 FLA 46.51 48.14 38.49 67.56 17.51 39.38 21.02 26.28 14.20 44.63 21.32 26.39 11.88 59.87 13.92 8.50 4.58 24.57 48.09 OMEN 23.43 24.94 11.03 55.07 13.13 15.97 4.21 6.69 2.92 18.8 4.13 6.72 3.12 47.41 2.53 0.63 0.99 4.48 23.41 PassFlow 10.38 11.35 1.41 27.92 5.34 4.82 0.70 0.30 0.05 5.01 0.41 2.29 0.21 24.02 1.79 0.45 0.16 0.39 9.54 PassGAN 6.94 7.65 0.16 28.31 0.30 3.55 0.10 0.23 0.08 4.12 0.10 0.53 0.12 24.17 0.10 0.01 0.07 0.06 7.53 PassGPT 38.47 39.66 33.74 46.55 25.82 29.81 20.04 19.90 11.19 33.43 17.88 29.62 10.39 41.68 18.19 12.72 5.04 21.63 36.16 PCFG 29.79 28.69 45.59 13.69 19.81 42.94 22.16 16.41 12.11 48.87 21.5 45.14 12.31 13.46 30.14 8.78 5.35 22.19 43.64 PLR-GAN 12.60 13.74 1.35 39.95 0.67 7.31 0.77 0.74 0.27 8.39 0.64 2.14 0.47 34.17 0.35 0.03 0.15 1.02 12.49 VGPT2 12.09 13.11 2.38 29.39 2.59 5.37 1.22 0.90 0.33 6.10 0.95 2.15 0.53 25.19 0.61 0.04 0.09 1.07 9.74
@@ 5.5.2 Analysis by Password Length
We analyzed matches by password length, ranging from 4 to 12 characters. Results, computed using a weighted average across all datasets, are shown in Figure 6. All models, except OMEN, perform well on short passwords, with FLA nearly achieving a $100\%$ match rate. However, performance declines as length increases. Beyond 8 characters, only PCFG, OMEN (up to 10 characters), FLA, and PassGPT maintain a substantial guessing rate. From length 9 onward, PCFG emerges as the most performing approach, surpassing both FLA and PassGPT, while, consistent with RQ1, PassGPT overtakes FLA from length 11. PassFlow initially performs on par with PassGPT for shorter lengths but undergoes a sharp decline, becoming the weakest performer from length 8 onward.
Figure 6: Guessing performance by password length. Dots show the percentage of generated passwords for each length.
@@ 5.5.3 Analysis by Password Patterns
We evaluated the models’ ability to guess passwords based on the structural patterns defined in Table IV. The results, presented in Table IX as a weighted average across all datasets, show that PassGPT, FLA, PCFG, and, in part, OMEN consistently achieve the strongest performance, standing out as the only models capable of successfully guessing passwords across all defined patterns. In contrast, the remaining models show limited pattern coverage, particularly struggling with complex password structures that include special characters (r5, r7, r8) or uncommon character combinations (r11, r12, r13). GAN-based models fare especially poorly on passwords composed entirely of special characters (r5), significantly underperforming PassFlow and VGPT2. This suggests a clear limitation in capturing rare or unconventional character sets. In contrast, all models exhibit strong performance on digit-only passwords (r4), indicating that simple patterns are consistently learned regardless of the architecture.
@@
We assess the extent to which approaches’ learned distributions overlap and investigate the effectiveness of combining multiple models to enhance guessing performance.
@@
We investigate the first aspect of this RQ utilizing two primary metrics: the Jaccard Index and the Mergeability Index. Their mathematical definitions are provided in Appendix B, in Equations 1 and 2, respectively. The Jaccard Index provides a quantitative measure of the overlap between two sets of generated passwords, with a value closer to 1 indicating that the two models produce similar distributions, and a value closer to 0 indicating that they generate distinct sets of passwords. As illustrated in Figure 7(a), FLA-PassGPT, FLA-PCFG, and FLA-OMEN pairs are the only combination with a Jaccard Index greater than $0.1$, suggesting that FLA shares some similarities with these models in password generation patterns. All other pairs exhibit much lower Jaccard values. Interestingly, despite PassGAN and PLR-GAN both being based on GANs, their Jaccard Index is quite low. Similarly, VGPT2-PassGPT also shows a low Jaccard Index. These results highlight that, even within the same architectural family, models may produce highly distinct password distributions, suggesting that combining multiple models could be beneficial for maximizing coverage of the password space.
@@
While the Jaccard Index provides insights into the overlap of generated passwords, it does not account for their effectiveness in terms of actual matches with real-world data. To address this, we introduce the Mergeability Index, a complementary metric that measures the benefits of combining the output of two models based on their successful guesses. It captures the performance improvement achieved by merging the guesses relative to the best-performing model. A Mergeability Index close to 0 indicates that the models guess mostly the same set of passwords, while a value close to 1 indicates largely distinct matching password sets. The results are shown in Figure 7(b). Interestingly, FLA exhibits the lowest mergeability across all models, indicating that its guessed passwords are generally a superset of those generated by the other models and that it has already captured a significant portion of the password space. Indeed, the only models that provide a non-negligible improvement to FLA are PCFG and, to a lesser extent, PassGPT and OMEN. In contrast, the other seven models show high mergeability between them, with their varying combinations improving successful guesses by up to 30%. This suggests that these models guess highly distinct sets of passwords, meaning they each capture different aspects of the password space.
@@
We assess the potential of a multi-model attack by evaluating the performance gain achieved through the combination of the generated password sets from $n$ models. Our results, presented in Figure 8, show the relative gain with respect to the best-performing single model, computed as percentage points on the test set. We identify the following optimal model combinations for different values of $n$: $n_{1}=\{FLA\}$, $n_{2}=n_{1}+PCFG$, $n_{3}=n_{2}+PassGPT$, $n_{4}=n_{3}+OMEN$, $n_{5}=n_{4}+PLR$-GANGAN, $n_{6}=n_{5}+PassFlow$, $n_{7}=n_{6}+VGPT2$, and $n_{8}=n_{7}+PassGAN$. In some datasets, combining models yields significant gains over FLA alone, around 12% for RockYou, and around 9% for LinkedIn and Ashley Madison, mainly thanks to PCFG and PassGPT. Remarkably, combining PCFG and FLA provides a larger average improvement than adding the remaining six models combined. These results mirror those presented in Figure 7(b), indicating that FLA, PCFG, and PassGPT capture somewhat complementary parts of the password space. In datasets like 000webhost and Mailru, FLA and PCFG merged perform nearly as well as the combination of all models, with less than a $0.5$ percentage point improvement when adding the other seven models.
@@ 5.7.1 Metrics-Based Evaluation
IMD is an intrinsic, multi-scale metric that compares the data manifolds of real and generated samples applicable across diverse domains. However, its effectiveness is highly sensitive to the choice of feature representation, which can introduce bias into the distance estimation.
@@
TABLE X: Distance between human- and generative model-created passwords.
@@
To contextualize the metric values, we define two baselines: (1) a soft lower bound (optimal case), defined as the metric value between test and training passwords; (2) a soft upper bound (worst case), defined as the metric value between randomly generated and test passwords. For each model, we normalize the metric value between lower ($0\%$) and upper ($100\%$) bounds, as displayed in Table X. Values close to zero indicate closer alignment between the generated and human-created password distributions. GAN-based models demonstrate generally good performance across all metrics, with PLR consistently achieving lower values than PassGAN. PassGPT generates passwords that closely resemble human-created ones, likely due to its underlying GPT2 architecture, which is well-suited for modeling textual distributions. OMEN’s passwords lie between human and random, achieving scores around 40-60% for most metrics. PCFG struggles only on IMD, with all other indicators tending towards human values. In contrast, PassFlow demonstrates the weakest performance, particularly in the IMD metric, where its score exceeds the soft upper bound by a factor of two. We attribute this behavior to PassFlow’s Gaussian Smoothing (GS) strategy, which introduces random perturbations in the generation process. While GS enhances uniqueness [50], it artificially distorts the distribution of generated passwords, ultimately reducing their quality. VGPT2 and FLA exhibit similar trends, performing strongly across most metrics except for IMD, where elevated values across most models suggest that they all struggle to fully capture certain characteristics of human passwords. Conversely, MTopDiv consistently yields low values, indicating that the models effectively avoid common mode failures, such as dropping, collapse, and invention.
@@ 5.7.2 Length Distribution in Generated Passwords
Figure 9 compares the length distribution of generated passwords with that of real passwords. Among the models, OMEN, VGPT2, FLA, PassFlow, and to a lesser extent PCFG, show the largest divergence from the real distribution: PassFlow tends to generate shorter passwords, FLA overestimates the probability of 8-character passwords, while OMEN, VGPT2, and PCFG of longer ones. In contrast, PassGAN, PLR-GAN, and PassGPT produce distributions that more closely align with real data, reinforcing previous observations.
@@ 5.7.3 Why Does IMD Yield High Values?
Figure 9 reveals a clear correlation between length distribution and IMD values. PassFlow, which exhibits one of the largest deviations from the real distribution, also records the highest IMD score. A similar trend is observed for FLA, VGPT2, OMEN, and PCFG. Conversely, models such as PassGAN, PLR-GAN, and PassGPT, which more closely replicate the real-world length distribution, achieve lower IMD scores. Interestingly, Table X shows that models such as PassFlow and FLA yield higher IMD scores than those associated with the random-password soft upper bound. We posit that this counterintuitive result stems from the uniform length distribution of random passwords. Although this distribution diverges significantly from that of real passwords, it nonetheless includes longer passwords that PassFlow and FLA struggle to generate. These observations suggest that IMD is highly sensitive to length distribution, favoring models that replicate it more accurately.
TABLE XI: Comparison summary of the selected approaches.
Metric FLA OMEN PFLOW PGAN PGPT PCFG PLR VGPT
@@
   Scenario: In-Distribution $\uparrow$ 49.21 29.07 12.76 8.85 37.21 37.97 14.88 11.79
   Scenario: Cross-Community $\uparrow$ 27.72 12.23 7.13 3.62 17.72 25.13 6.16 4.76
   Scenario: Cross-Culture $\uparrow$ 23.27 13.55 16.17 6.52 14.24 12.11 9.70 7.01
@@
   Frequency: Common (Top 5%) $\uparrow$ 83.37 66.65 31.20 27.87 68.20 77.39 36.86 32.83
   Frequency: Common (Top 10%) $\uparrow$ 76.60 57.64 25.73 22.73 63.45 67.50 30.63 27.23
   Frequency: Rare (Bottom 90%) $\uparrow$ 47.03 27.40 12.28 8.75 34.51 33.78 13.61 10.83
@@
   Length: Short (4-7 Chars) $\uparrow$ 72.97 32.42 28.16 19.91 55.62 49.60 32.96 29.65
   Length: Medium (8-10 Chars) $\uparrow$ 36.38 20.33 1.33 2.48 26.93 33.44 4.47 2.16
   Length: Long (11-12 Chars) $\uparrow$ 11.60 2.64 0.00 0.07 12.58 15.56 0.24 0.05
@@
   Pattern: Simple (1 Char Class) $\uparrow$ 53.52 33.97 16.22 14.06 41.16 24.42 21.71 17.85
   Pattern: Moderate (2 Char Classes) $\uparrow$ 38.94 15.69 4.71 3.46 29.56 42.40 7.15 5.26
   Pattern: Complex (3 Char Classes) $\uparrow$ 14.20 2.92 0.05 0.08 11.19 12.11 0.27 0.33
@@
   Train Set Size Sensitivity (%) $\downarrow$ 7.58 1.15 1.90 12.56 13.16 9.96 3.54 35.49
   Cross-Community Loss (%) $\downarrow$ 30.13 20.54 6.84 24.16 42.15 28.58 32.42 32.49
   Cross-Culture Loss (%) $\downarrow$ 58.53 62.13 14.58 48.67 66.15 69.31 49.27 57.11
@@
   Uniqueness (%) $\uparrow$ 100.00 100.00 90.10 54.59 73.22 99.89 71.13 91.41
   Humanness Distance (%) $\downarrow$ 38.50 40.33 70.16 19.83 2.40 19.16 4.80 44.00
5.8 Resource Consumption
Experiments were conducted on an Ubuntu 24.04.2 LTS server equipped with 2x AMD EPYC 9354 (32 cores per CPU, 2 threads per core), 503 GB of RAM, and 5x NVIDIA RTX A6000 with 48 GB of VRAM each. In Table XII we report resource consumption during training, computed over a training dataset of 1 million passwords with a maximum length of 12, and resource consumption during sampling, computed when generating $5\times 10^{8}$ passwords. Except for GANs, the reported training time multiplied by $x$ denotes the expected time to train on $x$ million passwords. By design, GAN-based approaches are independent of dataset size, as training stops after a fixed number of generator iterations; thus, the reported total training time reflects the expected time to train such models.
OMEN and PCFG exhibit clear advantages, as they are the fastest and most memory-efficient methods for training and sampling. Regarding DL-based approaches, the resource consumption highly depends on the model size, architecture, and guessing strategy. Although the training and sampling times are higher than those of traditional tools, they are not prohibitively so; the primary concern lies in memory usage. PassGPT, PassFlow, and partially FLA are computationally demanding, making them practical only in unconstrained environments (e.g., clusters). During sampling, FLA requires $\sim 28$ GB of RAM to complete the algorithm described in Appendix A. Similarly, PassFlow’s GS technique requires storing in memory all unique passwords generated for the whole sampling process, corresponding to $\sim 43$ GB of RAM when generating $5\times 10^{8}$ passwords. Instead, PassGPT requires at least a GPU equipped with >>24 GB of VRAM during training, and $\sim 51$ GB of RAM during sampling. The remaining three models: PassGAN, PLR-GAN, and VGPT2 are more resource-efficient, requiring at most 2 GB of RAM during training and 4 GB of RAM during sampling, making them suitable for deployment in various environments.
Regarding model size, for DL-based approaches, we report the size of the model itself. In contrast, for OMEN and PCFG, the reported size refers to the average disk space required to store the generated rules. FLA is by far the largest model among the others, weighing 391 MB. The transformer-based models, PassGPT and VGPT2, follow, with sizes slightly above 200 MB. All the other approaches can be considered lightweight, occupying less than 100 MB.
TABLE XII: Time and memory consumption of the selected approaches. Training stats follow the format h:m:s.
Training Stats Sampling Stats General Model 1 Epoch Total Time RAM Peak VRAM Peak PW/s RAM Peak VRAM Peak Model Size FLA 00:06:38 02:16:40 1.31GB 10.07GB 9926 28.74GB 4.97GB 391.2MB OMEN 00:00:39 00:00:39 0.07GB N/A 653594 0.77GB N/A 81.64MB PassFlow 00:00:45 02:30:00 1.95GB 0.47GB 17538 43.63GB 0.43GB 59.2MB PassGAN 00:05:58 12:43:44 2.13GB 0.43GB 77483 4.38GB 0.41GB 22.4MB PassGPT 00:06:52 00:19:36 2.65GB 23.63GB 3672 51.13GB 5.19G 230.2MB PCFG 00:00:40 00:00:40 0.07GB N/A 370370 0.67GB N/A 84.29MB PLR-GAN 00:09:00 38:24:00 2.16GB 0.40GB 62617 4.89GB 0.42GB 11.2MB VGPT2 00:03:26 01:01:20 1.90GB 1.31GB 4374 4.03GB 1.78GB 206.3MB